The firm's target capital structure is the mix of debt, preferred stock, and common equity the firm plans to raise funds for its future projects. The target proportions of debt, preferred stock, and common equity, along with the cost of these components, are used to calculate the firm's weighted average cost of capital (WACC). If the firm will not have to issue new common stock, then the cost of retained earnings is used in the firm's WACC calculation. However, if the firm will have to issue new common stock, the cost of new common stock should be used in the firm's WACC calculation.

Quantitative Problem: Barton Industries expects that its target capital structure for raising funds in the future for its capital budget will consist of 40% debt, 5% preferred stock, and 55% common equity. Note that the firm's marginal tax rate is 25%. Assume that the firm's cost of debt, rd, is 9.9%, the firm's cost of preferred stock, rp, is 9.1% and the firm's cost of equity is 12.5% for old equity, rs, and 12.9% for new equity, re. What is the firm's weighted average cost of capital (WACC1) if it uses retained earnings as its source of common equity? Do not round intermediate calculations. Round your answer to two decimal places.

  %

What is the firm’s weighted average cost of capital (WACC2) if it has to issue new common stock? Do not round intermediate calculations. Round your answer to two decimal places.

  %
Answer #1

After-tax cost of debt=9.9*(1-tax rate)

=9.9*(1-0.25)=7.425%

WACC=Respective costs*Respective weight

a.WACC=(7.425*0.4)+(9.1*0.05)+(12.5*0.55)

=10.3%

b.WACC=(7.425*0.4)+(9.1*0.05)+(12.9*0.55)

=10.52%